OpenStudy (anonymous):

Find the slope of the line that passes through the pair of points (17, 9) and (5, 29)

hero (hero):

1. insert points (17,9) and (5,29) into y = mx + b 29 = 5m + b 9 = 17m + b 2. Isolate b in each equation: b = 29 - 5m b = 9 - 17m 3. Set b = b 29 - 5m = 9 - 17m 4. Use inverse operations to place like terms on the same side: 29 - 9 = 5m - 17m 20 = -12m 5. Isolate m: 20/-12 = m -3/5 = m 6. Find b by inserting m into one of the equations in step 2: b = 29 - 5(-3/5) b = 29 + 3 b = 32 7. Insert the values of m and b into y = mx + b: y = (-3/5)x + 32 8. You're done

OpenStudy (ghazi):

just use \[m=\frac{ Y _{2}-Y _{1} }{ X _{2}-X _{1} }=\frac{ 29-9 }{ 5-17 }=\frac{ 20 }{ -12 }=\frac{ -5 }{ 3 }\]